Circular No. 7745

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE B. Barris, University of Hawaii, on behalf of the High-Z Supernova

Search Team (cf. IAUC 6160, 6646, 7312, 7516), reports the discovery of several apparent supernovae on I-band images taken with the Canada-France-Hawaii Telescope (+ 12K Mosaic in queue mode); the supernovae were identified by subtracting images taken on Sept. 11 from those taken Oct. 9, and confirmed on images taken on Sept. 24. Additional I-band confirmation images were obtained by P. Challis, B. Barris, P. Capak, and E. Hu (Subaru telescope + SUPRIME) on Oct. 16 and 22; by T. Hayward, K. Krisciunas, B. Schommer, R. C. Smith, and N. Suntzeff (Gemini South telescope + Acquisition Camera) on Oct. 20; and by A. Clocchiatti and R. C. Smith (European Southern Observatory 8.2-m Melipal UT3 Very Large Telescope + FORS1) on Oct. 21. Spectral observations were made at the VLT (Oct. 21), and also with the Keck II telescope (+ ESI) by A. Herrero and P. Najarro (Oct. 19, 20) and by A. V. Filippenko and R. Chornock (Oct. 22, 23). In the table below, I_{1}, I_{2}, and I_{3} indicate the I magnitudes on Oct. 9, 21, and 22, respectively. Each apparent supernova is within 2′′ of its host galaxy’s center.

SN  α_{2000}  δ_{2000}  I_{1}  I_{2}  I_{3}  z
2001fi  2^{h}26^{m}38.^{s}84 + 0^{o}26′15′′.3 22.7  23.0 0.32
2001fj  2 27 03.88 + 0 23 52.6 23.2  23.3 0.22
2001fk 2 28 04.67 + 0 46 18.0 23.2  23.7  ?
2001fl  2 28 15.18 + 0 21 18.6 20.2  19.6 0.087
2001fm 2 28 24.23 + 0 43 54.0 23.9 24.4  0.801
2001fn 2 29 00.47 + 0 42 21.8 22.3  22.9 0.186
2001fo 4 37 41.45 − 1 29 33.1 24.0 22.7  0.772
2001fp 4 38 04.11 − 1 42 32.1 24.0  24.3 0.76
2001fq 4 38 04.49 − 1 23 48.9 24.3 24.1  0.76?
2001fr 4 38 50.06 − 1 22 33.3 23.4  22.5 0.18
2001fs 4 39 30.68 − 1 28 21.9 24.4 23.9  0.84
2001ft  4 39 32.22 − 1 22 47.8 25.1 24.9  1.3

SN 2001fk was present on the Sept. 11 images but was not present on images taken in Oct./Nov. 1999 (cf. IAUC 7312). The remaining eleven apparent supernovae are invisible in the Sept. 11 images (limiting mag I = 25.8). SN 2001fl shows a type-IIn spectrum like that of SN 1988Z. SN 2001fs is a type-Ia supernova.

2001 November 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green